-4/9(2x-4)=48 please show me stepwise
Answer by kikerkins(17)

-4/9(2x-4)=48
Distribute the -4/9 to the 2X and -4
-8/9x+16/9=48
Then becuase we need the variables on one side we subtract the 16/9 from 48
-8/9= 42 2/9
Now we have -8/9X=380/9( Made it an improper fraction)
Now divide 380/9 by -8/9
You get X=-37 1/2